1. Dr. Phil wishes to know how many shares are necessary to elect 3 directors out...

1.

Dr. Phil wishes to know how many shares are necessary to elect 3 directors out of 9 directors up for election for the board of the Winfrey Publishing Company. There are 380,000 shares outstanding. (Do not round intermediate calculations.)
  

Number of shares

2.

Rust Pipe Co. was established in 1994. Four years later the company went public. At that time, Robert Rust, the original owner, decided to establish two classes of stock. The first represents Class A founders' stock and is entitled to ten votes per share. The normally traded common stock, designated as Class B, is entitled to one vote per share. In late 2010, Mr. Stone, an investor, was considering purchasing shares in Rust Pipe Co. While he knew the existence of founders’ shares were not often present in other companies, he decided to buy the shares anyway because of a new technology Rust Pipe had developed to improve the flow of liquids through pipes.

Of the 1,500,000 total shares currently outstanding, the original founder's family owns 51,925 shares.


What is the percentage of the founder's family votes to Class B votes? (Do not round intermediate calculations. Input your answer as a percent rounded to 2 decimal places.)

Percentage of votes %
Answer #1

1. No.of shares = (( Number of Directors Needed to be Elected x Outstanding Shares of the company) / (Number of Directors to be Elected + 1)) + 1

= ( (3 * 380,000) / ( 9 + 1) ) + 1

=  ( 1,140,000 / 10) + 1

= 114,000 + 1

= 114,001 shares.

2.The founder's family members holds Class A Founders Stock and are entitled to 10 votes per share.

No. of founder's family votes = 51,925 * 10 = 519,250 votes.

Shareholders holding Class B shares are entitled to 1 vote per share.

No.of Class B shareholders Votes = (1,500,000 - 51,925) * 1 = 1,148,075 votes

  The percentage of the founder's family votes to Class B votes = (519,250 / 1,148,075) * 100 = 45.23%

The percentage of the founder's family votes to Class B votes = 45.23%